OpenStudy (anonymous):

The point (2,n) lies on the circle whose equation is (x-3)^2 + (y+2)^2=6. Find the value of n. THANK YOU :)

OpenStudy (anonymous):

put \(x=2\) get \[(2-3)^2+(y+2)^2=6\] and solve for \(y\)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

But we need to find n not y

OpenStudy (anonymous):

lol it is a number makes no difference what variable you use right?

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

if you are married to the \(n\) put \[(2-3)^2+(n+2)^2=6\] and solve for \(n\)

OpenStudy (anonymous):

Haha

OpenStudy (anonymous):

Ill try that

OpenStudy (anonymous):

ok btw you should get two answers, because the circle will have to points on it with the first coordinate of \(2\)

OpenStudy (anonymous):

I got the square root of 5 and then subtract 2

OpenStudy (anonymous):

yes although it is really \[n+2=\pm\sqrt5\]

OpenStudy (anonymous):

because as i said there should be two answers

OpenStudy (anonymous):

So how exactly should I right the answer?

OpenStudy (anonymous):

Write*

OpenStudy (anonymous):



OpenStudy (anonymous):

\[n=-2\pm\sqrt5\] is how i would write it

OpenStudy (anonymous):

Ohh, ok thanks :D

OpenStudy (anonymous):

Wait is it -5 or +5?

OpenStudy (irishboy123):

Ashely, if you "got the square root of 5 and then subtract 2", there could be something amiss in your methodology and you might want to go through it again more slowly. if it helps, post your work here in your own time and someone will be here to help if that is needed.